I can't do word problems please help if you can. Thank you
20.)An annuity has a record of 9.58% annual growth compounded annually. how long will it take to double your money in this fund?

Answer by stanbon(75887)

An annuity has a record of 9.58% annual growth compounded annually. how long will it take to double your money in this fund?
---------
2A = A(1+r/n)^(nt)
2A = A (1+0.0958/1)^(t)
2=(1.0958)^t
Take the log of both sides to get;
log2 = t (log 1.0958)
t= 7.58 years to double your money
Cheers,
Stan H.